A racing car accelerates uniformly through 3 gear changes with the following average speeds: 20m/s for 2s,40m/s for 2s,60m/s for 6s.What is the overall average speed?
1
Expert's answer
2013-02-28T09:32:50-0500
V - an overal average speed
S - a total distance
t - a total time
V = S/t
S = V1*t1 + V2*t2 + V3*t3
V1, V2, V3 - average speeds for each gear changes
t = t1 + t2 + t3
t1, t2, t3 - times for each gear changes
Finally:
V = (V1*t1 + V2*t2 + V3*t3)/( t1 + t2 + t3) = (20*2+40*2+60*6)/(2+2+6)=48 m/s
Answer: the overall average speed equals 48 m/s